Derive an expression for Electric field due uniformly charged solid sphere at a) outside point b) inside point using Gauss's law
1
Expert's answer
2021-07-08T13:23:12-0400

Let, R be the radius of the shell and Q be the charge uniformly distributed on the surface. i) For a point  outside the shell :By Gauss's law,​"E.dA= \frac{Q_{en}}{\epsilon_0}"

"E.(4\pi r^2)=\frac{Q_{en}}{\epsilon_0}"

here r be the distance from centre of shell (r>R) and charge enclosed by surface "Q_{en}=Q" ​"E=\frac{kQ}{r^2}" So, Eout​"E_{out}=\frac{Q}{4\pi\epsilon_0 r^2}" ii) For a point inside the shell :By Gauss's law,

"E.A=\frac{Q_{en}}{\epsilon_0}"

​"E.(4\pi r^2)=\frac{Q_{en}}{\epsilon_0}" ​here r be the distance from centre of shell (r<R) and charge inside the shell, "Q_{en}=0" So 

"E.dA=0"


"E_{in}=0"